OpenStudy (anonymous):

A box contains 95 pink rubber bands and 90 brown rubber bands. You select a rubber band at random from the box. Find each probability. Write the probability as a fraction in simplest form. a. Find the theoretical probability of selecting a pink rubber band. b. Find the theoretical probability of selecting a brown rubber band.

OpenStudy (lucaz):

P(pink)=95/185 (simplify..) P(brown)=90/185 (simplify again)